HOW MUCH VOUME. OF WATER SHOULD BE ADDED TO CONVERT  100 ml of .1M SOLUTION INTO 0.01M SOLUTION

Grade:

3 Answers

vikas askiitian expert
13 years ago

initial volume of sol = 0.1

moles/V = M

moles = MV                   (M molality & V volume in liters)

moles of solute = 0.1*0.01 = .01

let volume of water added = V litre then final vol will be (0.1+V)

finall molarity = moles/volume = 0.01/(0.1+V) , this should be equal to 0.01

       0.001/(0.1+V) = 0.01

           V = 0.9L=900ml        ans

approve if u like it

sai raghava rajeev penagamuri
13 years ago

900ML

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Use molarity=moles/volume of the solution

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi